# AdvLoRA: Adversarial Low-Rank Adaptation of Vision-Language Models

### v1

@@ Abstract.
Vision-Language Models (VLMs) are a significant technique for Artificial General Intelligence (AGI). With the fast growth of AGI, the security problem become one of the most important challenges for VLMs. In this paper, through extensive experiments, we demonstrate the vulnerability of the conventional adaptation methods for VLMs, which may bring significant security risks. In addition, as the size of the VLMs increases, performing conventional adversarial adaptation techniques on VLMs results in high computational costs. To solve these problems, we propose a parameter-efficient Adversarial adaptation method named AdvLoRA by Low-Rank Adaptation. At first, we investigate and reveal the intrinsic low-rank property during the adversarial adaptation for VLMs. Different from LoRA, we improve the efficiency and robustness of adversarial adaptation by designing a novel reparameterizing method based on parameter clustering and parameter alignment. In addition, an adaptive parameter update strategy is proposed to further improve the robustness. By these settings, our proposed AdvLoRA alleviates the model security and high resource waste problems. Extensive experiments demonstrate the effectiveness and efficiency of the AdvLoRA.
@@
ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial General Intelligence (AGI), which aims to create intelligent agents that can perform as well as or better than humans on a wide range of cognitive tasks, is a promising topic for both research and industrial products (Pei et al., 2019; Goertzel, 2014). As vision and language are the most important information of intelligence, Vision-Language Models (VLMs) have become a significant technique for achieving AGI (Fei et al., 2022; Achiam et al., 2023).
In recent years, the adaptation of VLMs aims to improve the performance on different downstream tasks and has become a hot research topic. However, through extensive experiments, we find the vulnerability of the conventional adaptation methods, e.g., Full Fine-Tuning (FFT) (Wang et al., 2017; Wu et al., 2022; Zhang et al., 2022b), Linear Probe (LP), LoRA (Hu et al., 2021), Unidapter (Lu et al., 2023), and Aurora (Wang et al., 2023) for VLMs, which may bring significant security threats in various domains, such as facial recognition (Venkatesaramani et al., 2021; Sharif et al., 2016), medical analysis (Finlayson et al., 2019; Ma et al., 2021) and autonomous driving (Zhang et al., 2022a; Feng et al., 2021). As shown in Figure 1, we conduct adaptation experiments of VLMs on the natural and attacked data of the MSCOCO (Lin et al., 2014) and MSR-VTT (Xu et al., 2016) datasets. From these experimental results, we find that the average performance drops about 30.98% on the attacked data. To solve this problem, various techniques are proposed against adversarial attacks by data augmentation (Volpi et al., 2018; Morris et al., 2020), attack detection (Metzen et al., 2016; Liu et al., 2018) and adversarial training (Goodfellow et al., 2014; Liu et al., 2020). As the most effective defense strategy, adversarial training enhances the adversarial robustness of VLMs by retraining the model on mined adversarial examples (Madry et al., 2018; Szegedy et al., 2013; Pang et al., 2020).
@@
Figure 1. The vulnerability of vision-language model adaptation methods on natural data and attacked data of two datasets.
However, as the sizes of VLMs increase, the conventional adversarial training method with full parameter updating to improve the adversarial robustness of VLMs will lead to high computing and storage costs. In recent years, Parameter-Efficient Fine-Tuning (PEFT) technology has garnered widespread attention as a novel adaptation paradigm due to its significant success in adapting large-scale pre-trained models. PEFT technologies can adapt VLMs with extremely small additional tunable parameters and achieve comparable or better performance than FFT methods. While PEFT technologies have demonstrated remarkable success in natural scenarios, their application in adversarial attack scenarios remains largely uncharted territory. But simply applying the adversarial training on the conventional adaptation methods will lead to 1) limited defense performance and 2) high computational and storage costs. To verify our points, we visualize the adversarial robustness performance and the tunable parameter number of different adversarial adaptation methods in Figure 2. From the results, we find that the existing adaptation methods such as FFT and UniAdapter will lead to large parameter costs. Besides, LoRA, LP, and Aurora are not robust to adversarial attacks.
@@
To solve these problems, we aim to develop a parameter-efficient adversarial adaptation method termed AdvLoRA to effectively and efficiently improve the robustness of VLMs against attacks. At first, similar to LoRA, the intrinsic low-rank property of adversarial adaptation for VLMs is revealed. Secondly, we improve LoRA with a novel reparameterizing technology. Concretely, we regard the rank of LoRA as the number of cluster centers and utilize the clustering algorithm to reparameterize LoRA from the weight matrices of VLMs. The weight matrices are decoupled into the clustering centers and the clustering distribution matrices. Subsequently, we impose constraints on their product to align with the parameter distribution of the original weight matrix. Moreover, we design an adaptive parameter update strategy to improve the robustness further. Through these settings, we effectively and efficiently facilitate the adversarial adaptation of VLMs. The contributions of this paper are summarized as follows.
@@
  • We investigate and reveal the intrinsic low-rank property during the adversarial adaptation for vision-language models.
@@
  • We are the first to introduce the adversarial adaptation for vision-language models. Extensive experiments demonstrate the effectiveness and efficiency of our proposed method.
@@ 2. Related Work
2.1. Vision-Language Models
Vision-Language Models (VLMs) have demonstrated success in addressing diverse vision-language downstream tasks, including cross-modal retrieval(Zeng and Mao, 2022; Huang et al., 2023; Geigle et al., 2022) and cross-modal generation(Ramesh et al., 2021, 2022; Bao et al., 2023; Rombach et al., 2022). Leveraging large-scale multi-modal data and self-supervised training, VLMs learn semantic associations across modalities and establish a generalized multi-modal representation subspace, so that it can be regarded as fundamental models to address cross-modal tasks. The architecture of VLMs typically consists of three parts, the text encoder, the image encoder, and the multi-modal fusion module. The text encoder encodes textual data, typically utilizing structures such as word embeddings(Mikolov et al., 2013; Kim et al., 2021), BERT(Devlin et al., 2018; Lu et al., 2019), and GPT(Radford et al., 2018, 2021). The image encoder, on the other hand, encodes visual data, commonly employing structures like CNN(LeCun et al., 1998; Huang et al., 2020), Fast R-CNN(Girshick, 2015; Lu et al., 2019; Chen et al., 2020), and ViT(Kim et al., 2021; Li et al., 2021b, 2022). The multi-modal fusion module is typically based on the transformer architecture(Vaswani et al., 2017) and can be categorized into three forms: encoder-based(Lu et al., 2019; Radford et al., 2021; Li et al., 2021b), decoder-based(Ramesh et al., 2021), and encoder-decoder-based(Wang et al., 2021). Benefiting from the advantages of both the encoder and decoder architecture, encoder-decoder-based VLMs can achieve better performance on both cross-modal understanding and generating tasks, like BLIP(Li et al., 2022). In this paper, we explore the adversarial robustness of BLIP. Recently, with the success of Large Language Models (LLMs), researchers have embarked on the exploration of how to enhance VLMs’ comprehension of multimodal data through the integration of LLMs. Examples include BLIP-2 (Li et al., 2023), Flamingo (Alayrac et al., 2022), LLaVA (Liu et al., 2024b, 2023a), Qwen-VL (Bai et al., 2023), and others.
2.2. Parameter-efficient Tuning on VLMs
Parameter-efficient Tuning technologies are first proposed to alleviate the heavy training and storage cost in the process of adaptation of pre-trained models. Adaptation is an effective way to help pre-trained models solve the downstream tasks, which usually occurs when pre-trained models perform a specific task that task-related data distribution is different from training data. However, as the size of VLMs increases, traditional adaptation technologies such as FFT are inefficient and costly(Wang et al., 2017; Wu et al., 2022; Zhang et al., 2022b). Recently, inspired by methods from natural language processing(Houlsby et al., 2019; Hu et al., 2021; Li and Liang, 2021; Liu et al., 2023e; Zhang et al., 2023; Dettmers et al., 2023) and computer vision(Rebuffi et al., 2017; Jia et al., 2022; Bahng et al., 2022) domains, some approaches designed for VLM are proposed, which aim to adapt frozen VLMs to downstream tasks by introducing extremely small learnable parameters. Although it has fewer learnable parameters, its effect can equal or even exceed that of the full-parameters tuning. These can be toughly divided into three types: adapter-based(Lu et al., 2023; Gao et al., 2023), prompt-based(Zhou et al., 2022; Lu et al., 2022; Xing et al., 2022), and LoRA-based(Dai et al., 2023; Dou et al., 2023; Hayou et al., 2024; Meng et al., 2024; Zhong et al., 2024; Qiang et al., 2024; Liu et al., 2024a; Wang et al., 2024; Agiza et al., 2024; Zhao et al., 2024; Pan et al., 2024). LoRA-based approaches have received considerable attention due to their fewer learnable parameters, no additional input, and no additional inference latency. In this paper, we explore the LoRA-based approach to VLMs from the clustering perspective.
2.3. Adversarial Robustness on VLMs
Some researchers have proven that artificial neural networks including VLMs are vulnerable to human-unrecognized attacks (Li et al., 2020a; Cai et al., 2023; Zhao et al., 2023). Specifically, adding additional perturbations to input can cause VLMs to make the incorrect decision with high confidence. To improve adversarial robustness on VLMs, most works focus on data augmentation (Cai et al., 2023; Wortsman et al., 2022) and adversarial training (Gan et al., 2020; Mao et al., 2022). Considered one of the most effective methods, adversarial training can improve the adversarial robustness of VLMs by injecting adversarial inputs into the training procedure through a min-max formulation (Gan et al., 2020; Mao et al., 2022). However, adopting adversarial training to VLMs is costly due to the huge parameters that need to be updated. To our best knowledge, TeCoA (Mao et al., 2022) is so far the only adversarial training method using parameter-efficient tuning technology to improve the adversarial robustness of VLMs. However, TeCoA focuses on single-modality downstream tasks (image classification) and has not been studied in the cross-modal tasks. In this paper, we explore the adversarial training to LoRA-based approach and try to improve the adversarial robustness of cross-modal tasks with VLMs inexpensively.
2.4. Clustering Algorithm
Clustering is a fundamental yet challenging task that aims to group the samples into separated clusters in an unsupervised manner. Benefit from the ability to mine unlabeled data. At the early stage, various traditional clustering methods (Hartigan and Wong, 1979; Von Luxburg, 2007; Reynolds, 2009; Ester et al., 1996; Rodriguez and Laio, 2014) are proposed. For example, the classical $k$-Means clustering (Hartigan and Wong, 1979) group samples via iteratively updating the cluster centers and cluster assignment. In recent years, inspired by the stunning performance of deep learning, deep clustering (Li et al., 2021a; Ronen et al., 2022; Min et al., 2018; Asano et al., 2019; Qian et al., 2022; Li et al., 2020b) has become a fast-growing research spot. For example, Xie et. al. propose DEC (Xie et al., 2016) to perform clustering via deep learning. Concretely, they first initialize the cluster centers via $k$-Means clustering on samples and then optimize clustering distribution with a Kullback–Leibler divergence clustering loss (Xie et al., 2016). Additionally, IDEC (Guo et al., 2017) is proposed to improve DEC by reconstructing original information from latent embeddings. Besides, JULE (Yang et al., 2016) is proposed to perform clustering via iteratively learning data embeddings and clustering assignments. Similarly, DeepCluster (Caron et al., 2018) also updates the deep network according to the clustering assignments in turn. In addition, an online method named SwAV (Caron et al., 2020) clusters the data and keeps the consistency between cluster assignments produced from different views of the same image. In DINO (Caron et al., 2021), a momentum encoder is adopted to alleviate representation collapse. Furthermore, Qian proposes SeCu (Qian, 2023) by designing a stable cluster discrimination task and a new hardness-aware clustering criterion. In addition to image data, deep clustering is also widely applied to graphs (Liu et al., 2022a, d, 2023d, 2023c, c, 2023b), texts (Aggarwal and Zhai, 2012; Liu et al., 2022b; Jacksi et al., 2020; Saeed et al., 2020), and recommendation (Liu et al., 2024c). However, parameter clustering is relatively rare. It is worth taking advantage of the unsupervised learning ability of clustering to separate the parameters into different clusters.
@@ 3. Method
In Section 3.1, we first define the cross-modal retrieval. Subsequently, addressing the vulnerability of VLMs to adversarial attacks, we introduce an adversarial training module in Section 3.2 to enhance the model’s adversarial robustness. Finally, to mitigate the high cost associated with adversarial training, we present an adaptation module in Section 3.3, which maintains the VLMs’ adversarial robustness while reducing the expenses of adversarial training.
@@ 3.1. Task Definition
3.1.1. Cross-Modal Retrieval
Cross-modal retrieval aims to utilize information from one modality to retrieve semantically relevant information from another. We select cross-modal retrieval as our benchmark task due to its efficacy in assessing the quality of cross-modal representation learning in VLMs. Under adversarial attacks, cross-modal retrieval serves as an effective metric for evaluating whether models can learn robust feature representations.
Taking image-to-text retrieval as an example, given an image $v_{i}$, its semantic representation $\mathbf{z}_{i}^{v}=\mathcal{F}_{v}(v_{i})$ is used to compute the cosine similarity with each textual representation $\mathbf{z}_{j}^{w}$ within the text database as follows.
(1) $\text{sim}(\mathbf{z}_{i}^{v},\mathbf{z}_{j}^{w})=\frac{\mathbf{z}_{i}^{v}\cdot\mathbf{z}_{j}^{w}}{\|\mathbf{z}_{i}^{v}\|\|\mathbf{z}_{j}^{w}\|},$
where $\mathbf{z}_{j}^{w}=\mathcal{F}_{w}({w_{j}})$ represents the semantic representation derived from the textual data $w_{j}$ after feature extraction via the text encoder $\mathcal{F}_{w}$. Then we select the highest similarity text data as the retrieval results. Under adversarial attacked, robust VLMs could learn semantically invariant feature representations so that they will not be misled by small perturbations.
@@ 3.2. Adversarial Training Module
Extensive experimentation demonstrated that both VLMs and their variants adapted with PEFT methods are susceptible to adversarial attacks, as illustrated in Figure 1 and the Appendix. Consequently, in this subsection, we design an adversarial training module to enhance the adversarial robustness of VLMs. We begin by introducing the concept of adversarial attacks, followed by the presentation of adversarial training as an effective defense technology for enhancing adversarial robustness.
3.2.1. Adversarial Attack
Adversarial attacks $\delta$ is a tensor added to the natural image $v$, $v_{a}=v+\delta$, aiming to fool the model into making the incorrect decision as formulated.
@@ 3.2.1. Adversarial Attack
where $p$ donates the $p$-norm, and $\varepsilon$ donates the restriction value of values, which is often set to be smaller than $8/255$. Thus, the adversarial attacks are imperceptible to humans. In this paper, we focus on adversarial attacks on visual data, as attacks on natural language are readily perceptible to humans. Therefore, it is practically significant and more challenging to make attacks on visual data. Concretely, we utilize PGD (Madry et al., 2018) to generate $v_{a}$ as follows.
@@
3.2.2. Adversarial Training
Adversarial training technologies refer to retraining the model on attacked data, which can learn semantically invariant features under adversarial attacks. Adversarial training aims to minimize the following objective.

### v3

# Enhancing Adversarial Robustness of Vision-Language Models through Low-Rank Adaptation

@@ Abstract.
Vision-Language Models (VLMs) play a crucial role in the advancement of Artificial General Intelligence (AGI). As AGI rapidly evolves, addressing security concerns has emerged as one of the most significant challenges for VLMs. In this paper, we present extensive experiments that expose the vulnerabilities of conventional adaptation methods for VLMs, highlighting significant security risks. Moreover, as VLMs grow in size, the application of traditional adversarial adaptation techniques incurs substantial computational costs. To address these issues, we propose a parameter-efficient adversarial adaptation method called AdvLoRA based on Low-Rank Adaptation. We investigate and reveal the inherent low-rank properties involved in adversarial adaptation for VLMs. Different from LoRA, we enhance the efficiency and robustness of adversarial adaptation by introducing a novel reparameterization method that leverages parameter clustering and alignment. Additionally, we propose an adaptive parameter update strategy to further bolster robustness. These innovations enable our AdvLoRA to mitigate issues related to model security and resource wastage. Extensive experiments confirm the effectiveness and efficiency of AdvLoRA.
footnotetext: Equal Contribution.footnotetext: Corresponding Author.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial General Intelligence (AGI) aims to create intelligent agents that can perform as well as or better than humans on various cognitive tasks, making it a promising topic for research and industrial applications (Pei et al., 2019; Goertzel, 2014). As vision and language are key components of intelligence, Vision-Language Models (VLMs) have emerged as a crucial technique for achieving AGI (Fei et al., 2022; Achiam et al., 2023).
@@
Figure 1. Vulnerability of vision-language model adaptation methods to natural and adversarial data in MSCOCO (image-text data) (Lin et al., 2014) and MSR-VTT (video-text data) (Xu et al., 2016) datasets.
Recently, the adaptation of VLMs aims to improve the performance on different downstream tasks and has become a hot research topic. However, through extensive experiments, we find the vulnerability of the conventional adaptation methods, e.g., Full Fine-Tuning (FFT) (Wang et al., 2017; Wu et al., 2022; Zhang et al., 2022b), Linear Probe (LP), LoRA (Hu et al., 2021), Unidapter (Lu et al., 2023), and Aurora (Wang et al., 2023) for VLMs, which may bring significant security threats in various domains, such as facial recognition (Venkatesaramani et al., 2021; Sharif et al., 2016), medical analysis (Finlayson et al., 2019; Ma et al., 2021) and autonomous driving (Zhang et al., 2022a; Feng et al., 2021). As shown in Fig. 1, we conduct adaptation experiments of VLMs on the natural and attacked data of the MSCOCO (Lin et al., 2014) and MSR-VTT (Xu et al., 2016) datasets. From these experimental results, we find that the average performance drops about 30.98% on the attacked data. To solve this problem, various techniques are proposed against adversarial attacks by data augmentation (Volpi et al., 2018; Morris et al., 2020), attack detection (Metzen et al., 2017; Liu et al., 2018) and adversarial training (Goodfellow et al., 2014; Liu et al., 2020). As the most effective defense strategy, adversarial training enhances the adversarial robustness of VLMs by retraining the model on mined adversarial examples (Madry, 2017; Szegedy et al., 2013; Pang et al., 2020).
However, as the sizes of VLMs increase, the conventional adversarial training method with full parameter updating to improve the adversarial robustness of VLMs will lead to high computing and storage costs (Gan et al., 2020). In recent years, Parameter-Efficient Fine-Tuning (PEFT) technology has garnered widespread attention as a novel adaptation paradigm due to its significant success in adapting large-scale pre-trained models. PEFT technologies can adapt VLMs with extremely small additional tunable parameters and achieve comparable or better performance than FFT methods. While PEFT technologies have demonstrated remarkable success in natural scenarios, their application in adversarial attack scenarios remains largely uncharted territory. But simply applying the adversarial training on the conventional adaptation methods will lead to 1) limited defense performance and 2) high computational and storage costs. To verify our points, we visualize the adversarial robustness performance and the tunable parameter number of different adversarial adaptation methods in Fig. 2. From the results, we find that the existing adaptation methods such as FFT and UniAdapter will lead to large parameter costs. Besides, LoRA, LP, and Aurora are not robust to adversarial attacks.
@@
To solve these problems, we aim to develop a parameter-efficient adversarial adaptation method termed AdvLoRA to effectively and efficiently improve the robustness of VLMs against attacks. At first, similar to LoRA, the intrinsic low-rank property of adversarial adaptation for VLMs is revealed. Secondly, we improve LoRA with a novel reparameterizing technology. Concretely, we regard the rank of LoRA as the number of cluster centers and utilize the clustering algorithm to reparameterize LoRA from the weight matrices of VLMs. The weight matrices are decoupled into the clustering centers and the clustering distribution matrices. Subsequently, we impose constraints on their product to align with the parameter distribution of the original weight matrix. Moreover, we design an adaptive parameter update strategy to improve the robustness further. Through these settings, we effectively and efficiently facilitate the adversarial adaptation of VLMs. Our designs on low-rank for adversarial adaptation are motivated by the common dense direction theory (Allen-Zhu and Li, 2022), which demonstrates that low-rank adaptation in shallow convolutional neural networks are more suitable to effectively enhance their robustness. For the first time, this paper empirically verifies the applicability of this theory to VLMs and introduces a novel clustering-based initialization method for LoRA, facilitating the process of adversarial fine-tuning. The contributions of this paper are summarized as follows:
@@
  • We investigate and reveal the intrinsic low-rank property during the adversarial adaptation for VLMs.
@@
  • We are the first to introduce the adversarial adaptation for VLMs. Extensive experiments demonstrate the effectiveness and efficiency of our proposed method.
@@ 2. Related Work
Parameter-efficient Tuning on VLMs. Vision-Language Models (VLMs) excel in various vision-language tasks, such as cross-modal retrieval (Hao et al., 2023a, b; Hao and Zhang, 2024; Hao et al., 2021a, 2022, b) and generation (Bao et al., 2023; Rombach et al., 2022). However, they may struggle when task data diverges from training data, necessitating re-training or fine-tuning on task-specific datasets. Traditional adaptation methods like Full Fine-Tuning (FFT) become inefficient with larger VLMs, prompting the need for parameter-efficient tuning to reduce training and storage costs. Recent approaches inspired by natural language processing (Liu et al., 2023; Zhang et al., 2023a; Dettmers et al., 2023) and computer vision (Jia et al., 2022; Bahng et al., 2022) aim to adapt frozen VLMs using minimal tunable parameters, achieving results comparable to full parameter tuning. These methods fall into three categories: adapter-based (Lu et al., 2023; Gao et al., 2023), prompt-based (Zhou et al., 2022; Lu et al., 2022; Xing et al., 2022), and LoRA-based (Zhong et al., 2024; Qiang et al., 2024; Liu et al., 2024b; Wang et al., 2024; Zhao et al., 2024; Pan et al., 2024). LoRA-based approaches are particularly notable for their reduced tunable parameters, lack of additional input, and minimal inference latency. This paper addresses suboptimal initialization in standard LoRA methods and explores a clustering-based reparameterization strategy to enhance VLM robustness during adaptation.
Adversarial Adaptation on VLMs. Some researchers have shown that artificial neural networks, including Vision-Language Models (VLMs), are vulnerable to unrecognized attacks (Cai et al., 2023; Zhao et al., 2023; Liu et al., 2025, 2024a). Specifically, adding perturbations to inputs can lead VLMs to make incorrect decisions with high confidence. To enhance adversarial robustness, many studies focus on data augmentation (Cai et al., 2023; Wortsman et al., 2022) and adversarial training (Gan et al., 2020; Mao et al., 2022; Diao et al., 2024). Adversarial training, particularly effective, improves robustness by incorporating adversarial inputs into the training process via a min-max formulation (Madry, 2017). Initially, some efforts used adversarial training techniques to train VLMs from scratch (Gan et al., 2020). Recently, adversarial adaptation has emerged as a cost-effective strategy for enhancing adversarial robustness post-pretraining (Xu et al., [n. d.]; Mao et al., 2022; Yuan et al., 2024; Zhang et al., 2023b). However, most methods update all parameters of the pre-trained model and focus on visual models. A few multi-modal approaches, such as TeCoA (Mao et al., 2022), utilize prompt tuning for adversarial adaptation.
In this paper, we introduce a parameter-efficient LoRA method for adversarial adaptation in cross-modal tasks. Unlike AutoLoRa (Xu et al., [n. d.]), which updates all parameters while addressing gradient instability, our approach maintains efficiency. Inspired by dense direction theory (Allen-Zhu and Li, 2022), which suggests that low-rank adaptation enhances robustness in shallow networks, we provide the first empirical validation of this theory in VLMs and present a novel clustering-based initialization method for LoRA to streamline adversarial fine-tuning.
@@ 3. Method
In Sec. 3.1, we first define the cross-modal retrieval. Subsequently, addressing the vulnerability of VLMs to adversarial attacks, we introduce an adversarial training module in Sec. 3.2 to enhance the model’s adversarial robustness. Finally, to mitigate the high cost associated with adversarial training, we present an adaptation module in Sec. 3.3, which maintains the VLMs’ adversarial robustness while reducing the expenses of adversarial training.
@@ 3.1. Task Definition
Cross-Modal Retrieval. Cross-modal retrieval aims to utilize information from one modality to retrieve semantically relevant information from another. We select cross-modal retrieval as our benchmark task due to its efficacy in assessing the quality of cross-modal representation learning in VLMs. Under adversarial attacks, cross-modal retrieval serves as an effective metric for evaluating whether models can learn robust feature representations.
Taking image-to-text retrieval as an example, given an image $v_{i}$, its semantic representation $\boldsymbol{z}_{i}^{v}=\mathcal{F}_{v}(v_{i})$ is used to compute the cosine similarity with each textual representation $\boldsymbol{z}_{j}^{w}$ within the text database as follows:
(1) $\text{sim}(\boldsymbol{z}_{i}^{v},\boldsymbol{z}_{j}^{w})=\frac{\boldsymbol{z}_{i}^{v}\cdot\boldsymbol{z}_{j}^{w}}{\|\boldsymbol{z}_{i}^{v}\|\|\boldsymbol{z}_{j}^{w}\|},$
where $\boldsymbol{z}_{j}^{w}=\mathcal{F}_{w}({w_{j}})$ represents the semantic representation derived from the textual data $w_{j}$ after feature extraction via the text encoder $\mathcal{F}_{w}$. Then we select the highest similarity text data as the retrieval results. Under adversarial attacked, robust VLMs could learn semantically invariant feature representations so that they will not be misled by small perturbations.
@@ 3.2. Adversarial Training Module
Extensive experimentation demonstrated that both VLMs and their variants adapted with PEFT methods are susceptible to adversarial attacks, as illustrated in Fig. 1. Consequently, in this subsection, we design an adversarial training module to enhance the adversarial robustness of VLMs. We begin by introducing the concept of adversarial attacks, followed by the presentation of adversarial training as an effective defense technology for enhancing adversarial robustness.
Adversarial Attack. Adversarial attacks $\delta$ is a tensor added to the natural image $v$, $v_{a}=v+\delta$, aiming to fool the model into making the incorrect decision as formulated:
@@
where $p$ donates the $p$-norm, and $\varepsilon$ donates the restriction value of values, which is often set to be smaller than $8/255$. Thus, the adversarial attacks are imperceptible to humans. In this paper, we focus on adversarial attacks on visual data, as attacks on natural language are readily perceptible to humans. Therefore, it is practically significant and more challenging to make attacks on visual data. Concretely, we utilize PGD (Madry, 2017) to generate $v_{a}$ as follows:
@@
Adversarial Training. Adversarial training technologies refer to retraining the model on attacked data, which can learn semantically invariant features under adversarial attacks. Adversarial training aims to minimize the following objective: